# An Empirical Study of Self-Admitted Technical Debt in Machine Learning Software

### v2

@@ 3.2. RQ2: What are the different types of SATD in ML-based systems?
Bavota and Russo (bavota2016large, ) have developed a taxonomy of SATDs in traditional software systems. Given the specific nature of ML code as described by Sculley et al. (Sculley, ), new types of technical debt not mentioned in this taxonomy may exist. Therefore, in this research question, we investigate the composition of SATD in ML-based software, by qualitatively investigating a sample of SATD comments in ML software. Our goal is to uncover types of SATDs that may be specific to ML code and identify the types of SATDs that are prevalent in ML code. Such an understanding can ML practitioners help be vigilant about the types of debt that exist in ML software.
@@ 3.3. RQ3: Which stages of the ML pipeline are more prone to SATD?
A typical ML software pipeline (amershi2019software, ) involves stages for reading data and cleaning data: 1) transforming data into features that can be consumed by the ML model; 2) training the model; 3) evaluating and validating the model; 4) and finally deploying the model (amershi2019software, ). During the evolution of the system, engineering teams often have to update these stages to experiment with new models or adjust to changes in operation data distributions. Through this process, as observed by Sculley et al. (Sculley, ), different types of technical debt can be introduced. In this research question, we examine the proportion of SATD contained in different stages of the studied projects by taking a sample of SATD comments. By knowing which stages are the most prone to technical debt, engineering teams will be able to better prioritize their maintenance activities and better support their ML engineers, e.g., with best practices and–or efficient tools.
@@ 3.4. RQ4: How long does SATD survive in ML-based systems?
Several empirical studies (Sculley, ; bavota2016large, ) report that technical debt leads to low quality, particularly in terms of maintainability, and makes further changes more expensive in the long run (potdar2014exploratory, ; nord2012search, ; kruchten2012technical, ; Alves2016, ). Technical debt needs servicing, which may involve refactoring code, improving tests, and deleting dead code to enable future improvements and reduce errors. An understanding of the empirical estimates for such servicing of technical debt (TD) would help ML stakeholders better estimate project effort and timelines. For instance, the time taken for TD introduction indicates the fast-paced accumulation of issues like untested ML configurations, data mismanagement, or lack of robust ML practices. Conversely, the time taken to remove debt can be dependent on factors such as triaging debt based on its perceived urgency, developer availability, complexity in understanding and managing the ML process, or even breaking changes across other ML modules. Hence, by examining the timing of introduction and removal of SATD through survival analysis, and comparing these patterns between ML and non-ML projects, we aim to provide insights into how the unique characteristics of ML development impact the evolution of technical debt.
@@ 3.5. RQ5: What are the characteristics of long-lasting SATDs in ML-based systems?
Persistent long-term accumulation of technical debt often signals areas in the codebase suffering from insufficient maintenance (tom2013exploration, ). Debt lingering in the codebase for an extended period is crucial to understand since it could be overlooked or ignored by developers based on their perceived lack of its importance, or perhaps the project maintainers continually defer addressing this debt to subsequent maintenance cycles. To avoid such practices, it is essential to understand the factors leading to this long-term accumulation of debt. This understanding can help ML development teams acquire the necessary information to redirect their focus towards debt more likely to linger for a long time, thereby preventing its long-term accumulation.
@@ 4.1.1. Choice of ML Domains
Applications of ML algorithms (ML-based systems) are prevalent in Healthcare, Finance, Retail, Travel, Social Media, etc. (dargan2020survey, ; deng2016deep, ; kamath2019deep, ; singh2021applications, ). All such industrial applications can be grouped into three high-level domains, namely, Image Processing, Natural Language Processing, and Audio Processing (das2015applications, ). Recently, the use of reinforcement learning in Autonomous Gameplay has been gaining a lot of attention from academia and the industry (albrecht2018autonomous, ), prompting us to include it in our dataset as well. Additionally, we include the Self Driving Car domain due to its widespread use in industry and the significant attention it has received from researchers (badue2021self, ). The Self Driving Car domain mostly leverages computer vision in their perception modules and is an extended version of Image Processing (ma2020artificial, ). While it is impossible to include all industrial domains, we select these five as representative examples (pouyanfar2018survey, ) of successful ML implementations. Finally, to account for other domains, we also add raw ML implementations like maintained ML courses that we could not fit into the five selected domains, labeling them as “Other ML” (see Figure 2). Understanding technical debt in machine learning courses is important; many practitioners initiate their ML journey through such courses. It is possible that within these courses, comprehensive practices for managing technical debt might not be thoroughly covered. Further, by incorporating the management of technical debt from an early stage, learners can develop robust foundations, promoting best practices applicable in diverse, real-world scenarios.
@@ 4.1.2. Keyword-Based Searching to get Candidate Repositories
For our analysis of the above six domains, we gathered ML-based open source projects from GitHub [^3] Accessed on: 2024-03-22, https://github.com/. We used the GitHub Search API v3444Accessed on: 2024-03-22,https://docs.github.com/en/free-pro-team@latest/rest to look for keywords specific to each domain (see Table 1) and obtain a list of relevant projects. Our choice of keywords is aligned with the common academic and industry terminologies, appended to the term “machine learning” to enable the GitHub search engine focus on ML implementations. For instance, “autonomous gameplay” was initially used to obtain a broad list of gaming projects. Then, the updated keywords, “game reinforcement learning” and “game machine learning” were used to enable the GitHub search engine focus on the ML implementations. For other domains, the synonyms for the domain were used as well, e.g., “audio machine learning” and “sound machine learning”. Only for the NLP domain, we used an application of NLP, i.e. “chatbot” due to the popularity of chat-based NLP ML architectures, e.g., transformers on which GPT-based models are built (adamopoulou2020overview, ). Our rationale of using “chatbot” was based on the fact that some repositories may not specifically mention “nlp”, but could be using NLP-based ML via chatbots.
As suggested by prior studies ((kalliamvakou2014promises, ; mirhosseini2017can, ; coelho2020github, )), for each set of keywords, we identify the top matching projects by using the “best match” feature of GitHub search API. For each search query, we leveraged the GitHub search API to look for the top 50 repositories, then moved on to the other keywords for that domain.
@@ 4.1.3. Filtering of toy-projects
Based on previous studies (kalliamvakou2014promises, ), we filtered out “toy projects” from our comprehensive list of relevant projects for each domain, using a two level approach. At the lowest level, we employed a threshold on typical project-related metrics to filter out “toy-projects”, which is a fairly common practice data collection strategy in Software Engineering research. For example, to study feature toggles in python projects, Hoyos et al. (hoyos2021removal, ) used Github Search API v3 to obtain a list of relevant projects, then filtered them out based on programmatic and manual checks. Among many others, Dabic et al. (dabic2021sampling, ) used a threshold of 10 stars to qualify mature projects. Alfadel et al. (alfadel2021use, ) used top-most starred, non-forked projects with at least 20 commits for studying dependabot security pull requests, while Chen and Jiang (chen2020studying, ) selected non-forked projects having at least 30 stars for studying the rationale behind using logging practices in Java (e.g., internationalization of the log messages).
On top of this threshold-based filtering, our second level filtering further refines our sampled repositories using manual checks of the selected repositories based on a systematic approach recommended by Basili et al. (basili1986experimentation, ). This approach is widely adopted in empirical software research. For instance, to study the characteristics of AIOps projects, Aghili et al. (aghili2023studying, ) obtained their dataset by searching GitHub using the keyword “AiOps” and filtered the resulting repositories through manual verification to determine whether each repository actually focused on AIOps, a process akin to our research.
@@
    R1. Forks are copies of a given repository with minor variations. To avoid duplication and maintain originality, removing forked projects is necessary, as done by other research like (bhatia2023towards, ; dabic2021sampling, ).
@@
    The use case of the research determines the choice of metric on which a threshold is applied. For instance, Hoyos et al. (hoyos2021removal, ) employed thresholds on data size, requiring the selected projects to have a size of more than 1MB. Aghili et al. (aghili2023studying, ) used a threshold of #$\#stars>1$, while Dabic et al. (dabic2021sampling, ) used #$\#stars>10$.
    For our case, as indicated by Amershi et al. (amershi2019software, ), ML pipelines contain stages like data processing, model building, etc. Given our focus on mature ML projects, we heuristically determined that such projects should typically implement the different stages of Amershi’s ML pipeline, and do so in separate modules. Therefore, a project must include at least five Python files corresponding to these stages to be considered mature for our study.
@@
  • C4. A project should implement machine learning.
    R4. Despite our meticulously designed keywords containing the search keyword “machine learning” presented in Table 1, it may be possible that the GitHub search API was unable to find a relevant project implementing ML. Hence, this criterion is in place.
@@
    Top-level check: Each project’s README was reviewed. As identified by Prana et al. (prana2019categorizing, ), READMEs include sections on the project’s purpose, installation instructions, and usage examples, providing clear information about the project’s intent.
@@
    Intermediate check: If information regarding C4 and C5 was not discernible from the previous step, we proceeded to examine the files in the source code directory for the presence of ML system stages. Heuristically, the implementation of an ML system involves filenames being indicative of Amershi’s ML pipeline stages (e.g., model.py, train.py, or preprocess.py).
@@
    Code check: If the previous steps did not yield valuable information, we conducted a finer-grained examination, which entailed 1) checking for the presence of ML library imports (e.g., tensorflow, sklearn, keras) and 2) inspecting for ML implementation specifics. The rationale for checking ML imports is based on the fact that ML-based systems utilize these libraries to build ML models.
Table 1. Queries used to find projects from various ML domains and a non-ML domain.
Type Domain Queries Used
ML Self Driving Car “autonomous driving machine learning”, “self driving machine learning”
Audio Processing “sound machine learning”, “audio machine learning”, “voice machine learning”, “speech machine learning”
Autonomous Gameplay “game reinforcement learning”, “game machine learning”
Image Processing “image processing machine learning”, “vision machine learning”, “image processing deep learning”
Natural Language Processing “natural language processing machine learning”, “chatbot machine learning”
Other ML “machine learning”
@@
4.2. Procurement of Non-ML Projects
To answer the research questions RQ1 and RQ4 mentioned in the Introduction, we also need a set of non-ML projects as baseline. To allow a fair comparison we need the following: 1) same programming language as that of ML projects 2) same count of repositories as the total count of ML projects (i.e., 318) 3) different technology, but similar level of complexity that would encapsulate architectural decisions, configuration management challenges as that of ML
Hence, using a similar process as outlined in the previous Subsection 4.1, we collected 318 (same count as that of ML projects) non-ML Python projects from GitHub. For the keyword-based selection mechanism of ML projects, we used the keywords “python projects”, “server” and “” (empty keywords) to obtain a list of non-ML python projects sorted by the “best match” criterion of GitHub Search. The rationale for using “server” is based on the fact that this keyword indicates backend implementations, and is a non-specific term, i.e., the term server can have a diverse meaning in different software contexts, which makes it a non-intrusive term for finding a candidate non-ml repositories. Such projects often entail complex architectural decisions, configuration management challenges, and performance optimization issues, which are fertile grounds for the accumulation of technical debt.
@@ 4.4. SATD Detector Tool
As stated in our literature review Section 2.3, many SATD detector tools used to identify instances of SATD from code comments. For our study, we use the state of the art SATD detection tool [^5] Accessed on: 2024-03-22, https://github.com/Tbabm/SATDDetector-Core, proposed by Liu et al. (liu2018satd, ), which employs a NLP-based pre-trained text classification ML model to automatically detect the presence of SATD in a code comment. As compared to other contemporary SATD detection tools that rely on pattern matching, this tool demonstrates a higher recall (liu2018satd, ).
This tool’s use of an NLP-based model to determine whether a specific comment signifies SATD contrasts most of the other approaches, which merely scan for words like “todo” or “fixme” using string matching techniques. Therefore, the tool can successfully identify regular comments such as “300 iterations seems good enough but you can certainly train longer” or “naive approach for generating forward mapping this is naive and probably not robust” as SATD even when such strings lack an explicitly defined “TODO”. This SATD detector has been utilized in previous studies. For instance, Liu et al. (liu2021exploratory, ) used this tool for studying SATD in deep learning frameworks, while Zampetti et al. (Zampetti, ) used this tool for recommending when design technical debt should be self admitted.

### v3

@@ 3.2. RQ2: What are the different types of SATD in ML-based systems?
Bavota and Russo (bavota2016large) have developed a taxonomy of SATDs in traditional software systems. Given the specific nature of ML code as described by Sculley et al. (Sculley), new types of technical debt not mentioned in this taxonomy may exist. Therefore, in this research question, we investigate the composition of SATD in ML-based software, by qualitatively investigating a sample of SATD comments in ML software. Our goal is to uncover types of SATDs that may be specific to ML code and identify the types of SATDs that are prevalent in ML code. Such an understanding can ML practitioners help be vigilant about the types of debt that exist in ML software.
@@ 3.3. RQ3: Which stages of the ML pipeline are more prone to SATD?
A typical ML software pipeline (amershi2019software) involves stages for reading data and cleaning data: 1) transforming data into features that can be consumed by the ML model; 2) training the model; 3) evaluating and validating the model; 4) and finally deploying the model (amershi2019software). During the evolution of the system, engineering teams often have to update these stages to experiment with new models or adjust to changes in operation data distributions. Through this process, as observed by Sculley et al. (Sculley), different types of technical debt can be introduced. In this research question, we examine the proportion of SATD contained in different stages of the studied projects by taking a sample of SATD comments. By knowing which stages are the most prone to technical debt, engineering teams will be able to better prioritize their maintenance activities and better support their ML engineers, e.g., with best practices and–or efficient tools.
@@ 3.4. RQ4: How long does SATD survive in ML-based systems?
Several empirical studies (Sculley; bavota2016large) report that technical debt leads to low quality, particularly in terms of maintainability, and makes further changes more expensive in the long run (potdar2014exploratory; nord2012search; kruchten2012technical; Alves2016). Technical debt needs servicing, which may involve refactoring code, improving tests, and deleting dead code to enable future improvements and reduce errors. An understanding of the empirical estimates for such servicing of technical debt (TD) would help ML stakeholders better estimate project effort and timelines. For instance, the time taken for TD introduction indicates the fast-paced accumulation of issues like untested ML configurations, data mismanagement, or lack of robust ML practices. Conversely, the time taken to remove debt can be dependent on factors such as triaging debt based on its perceived urgency, developer availability, complexity in understanding and managing the ML process, or even breaking changes across other ML modules. Hence, by examining the timing of introduction and removal of SATD through survival analysis, and comparing these patterns between ML and non-ML projects, we aim to provide insights into how the unique characteristics of ML development impact the evolution of technical debt.
@@ 3.5. RQ5: What are the characteristics of long-lasting SATDs in ML-based systems?
Persistent long-term accumulation of technical debt often signals areas in the codebase suffering from insufficient maintenance (tom2013exploration). Debt lingering in the codebase for an extended period is crucial to understand since it could be overlooked or ignored by developers based on their perceived lack of its importance, or perhaps the project maintainers continually defer addressing this debt to subsequent maintenance cycles. To avoid such practices, it is essential to understand the factors leading to this long-term accumulation of debt. This understanding can help ML development teams acquire the necessary information to redirect their focus towards debt more likely to linger for a long time, thereby preventing its long-term accumulation.
@@ 4.1.1. Choice of ML Domains
Applications of ML algorithms (ML-based systems) are prevalent in Healthcare, Finance, Retail, Travel, Social Media, etc. (dargan2020survey; deng2016deep; kamath2019deep). All such industrial applications can be grouped into three high-level domains, namely, Image Processing, Natural Language Processing, and Audio Processing (das2015applications). Recently, the use of reinforcement learning in Autonomous Gameplay has been gaining a lot of attention from academia and the industry (albrecht2018autonomous), prompting us to include it in our dataset as well. Additionally, we include the Self Driving Car domain due to its widespread use in industry and the significant attention it has received from researchers (badue2021self). The Self Driving Car domain mostly leverages computer vision in their perception modules and is an extended version of Image Processing (ma2020artificial). While it is impossible to include all industrial domains, we select these five as representative examples (pouyanfar2018survey) of successful ML implementations. Finally, to account for other domains, we also add raw ML implementations like maintained ML courses that we could not fit into the five selected domains, labeling them as “Other ML” (see Figure 2). Such repositories are ML implementations with a lot of community activity. For instance, the project ML-Glossary [^3] Accessed on: 2025-04-23 https://github.com/bfortuner/ml-glossary has 3.1+K stars, 700+ forks, 135 watchers, and the Readme reports  54K monthly users—demonstrating strong community adoption. Understanding technical debt in machine learning courses is important; many practitioners initiate their ML journey through such courses, and the non-trivial adoption of these projects clearly represents their role in the ML open source ecosystem. It is possible that within these courses, comprehensive practices for managing technical debt might not be thoroughly covered. Further, by incorporating the management of technical debt from an early stage, learners can develop robust foundations, promoting best practices applicable in diverse, real-world scenarios.
Table 1. Queries used to find projects from various ML domains and a non‐ML domain, with rationale for synonym selection.
Domain Queries Used Rationale
Self Driving Car “autonomous driving machine learning”, “self driving machine learning” Reflects industry term (“autonomous driving”) vs common parlance (“self driving”) to capture variations in project naming.
Audio Processing “sound machine learning”, “audio machine learning”, “voice machine learning”, “speech machine learning” Covers raw acoustics (sound), general audio tasks, human speaker‐specific projects (voice and speech), ensuring broad coverage across subfields.
Autonomous Gameplay “autonomous gameplay machine learning”, “game machine learning”, “game reinforcement learning”, Captures varied terminologies (e.g. RL or ML) used by github projects for implementing ML in games
Image Processing “image processing machine learning”, “vision machine learning”, “image processing deep learning” Captures classical terminology (image processing), domain context (vision), and DL‐centric methods for modern implementations.
Natural Language Processing “natural language processing machine learning”, “chatbot machine learning” Includes explicit NLP label and application‐specific term (chatbot) to catch transformer‐based conversational projects that may omit “NLP” tag.
Other ML “machine learning” Serves as a catch‐all to include any ML repositories not covered by domain‐specific keywords. These were mostly educational projects.
@@ 4.1.2. Keyword-Based Searching to get Candidate Repositories
For our analysis of the above six domains, we gathered ML-based open source projects from GitHub [^4] Accessed on: 2025-04-23, https://github.com/. We used the GitHub Search API v3555Accessed on: 2025-04-23,https://docs.github.com/en/free-pro-team@latest/rest to look for keywords specific to each domain (see Table 1) and obtain a list of relevant projects. To ensure the relevance of a repository to ML, we appended each domain’s keywords term to “machine learning”. Given that the Github Search engine retrieval is based on keyword matching in the repository name, description or contents of the readme file [^6] Accessed on: 2025-04-23, https://docs.github.com/en/search-github/getting-started-with-searching-on-github/understanding-the-search-syntax, we used a comprehensive set of keywords for each domain to ensure a high recall (i.e., from all available repositories pertaining to a topic, we should not miss any) in our selection process. We also ensure high precision by employing a comprehensive filtering process explained in the subsequent subsections. While our selected keywords capture the generic and specific naming convention within each domain, detailed rationales for each domain‘s keywords are Table 1.
As suggested by prior studies (kalliamvakou2014promises; mirhosseini2017can; coelho2020github), for each set of keywords, we identify the top matching projects by using the “best match” feature of GitHub search API. For each search query, we leveraged the GitHub search API to look for the top 50 repositories, then moved on to the other keywords for that domain.
@@ 4.1.3. Filtering of toy-projects
Based on previous studies (kalliamvakou2014promises), we filtered out “toy projects” from our comprehensive list of relevant projects for each domain, using a two level approach. At the lowest level, we employed a threshold on typical project-related metrics to filter out “toy-projects”, which is a fairly common practice data collection strategy in Software Engineering research. For example, to study feature toggles in python projects, Hoyos et al. (hoyos2021removal) used Github Search API v3 to obtain a list of relevant projects, then filtered them out based on programmatic and manual checks. Among many others, Dabic et al. (dabic2021sampling) used a threshold of 10 stars to qualify matured projects. Alfadel et al. (alfadel2021use) used top-most starred, non-forked projects with at least 20 commits for studying dependabot security pull requests, while Chen and Jiang (chen2020studying) selected non-forked projects having at least 30 stars for studying the rationale behind using logging practices in Java (e.g., internationalization of the log messages).
On top of this threshold-based filtering, our second level filtering further refines our sampled repositories using manual checks of the selected repositories based on a systematic approach commonly used in empirical software research. For instance, to study the characteristics of AIOps projects, Aghili et al. (aghili2023studying) obtained their dataset by searching GitHub using the keyword “AiOps” and filtered the resulting repositories through manual verification to determine whether each repository actually focused on AIOps, a process akin to our research.
@@
    R1. Forks are copies of a given repository with minor variations. To avoid duplication and maintain originality, removing forked projects is necessary, as done by other research (bhatia2023towards; dabic2021sampling).
@@
    The use case of the research determines the choice of metric on which a threshold is applied. For instance, Hoyos et al. (hoyos2021removal) employed thresholds on data size, requiring the selected projects to have a size of more than 1MB. Aghili et al. (aghili2023studying) used a threshold of #stars>1\#stars>1, while Dabic et al. (dabic2021sampling) used #stars>10\#stars>10.
    For our case, as indicated by Amershi et al. (amershi2019software), ML pipelines contain stages like data processing, model building, etc. Given our focus on mature ML projects, we heuristically determined that such projects should typically implement the different stages of Amershi’s ML pipeline, and do so in separate modules. Therefore, a project must include at least five Python files corresponding to these stages to be considered mature for our study.
@@
  • C4. A project should implement a machine learning pipeline.
    R4. Our carefully designed search keywords (see Table 1) including the keyword “machine learning” should ideally enable the GitHub search API to retrieve ML‐related repositories. However, there may be noise in this process, i.e., a non‐ML repo might have been recommended by GitHub’s search engine. This is why criterion C4 is in place.
@@
    Top-level check: Each project’s README was reviewed. As identified by Prana et al. (prana2019categorizing), READMEs include sections on the project’s purpose, installation instructions, and usage examples, providing clear information about the project’s intent.
@@
    Intermediate check: Intermediate check: If information regarding C4 and C5 was not discernible from the previous step, we proceeded to examine the files in the directory for the presence of ML system stages. Heuristically, the implementation of an ML system involves filenames being indicative of Amershi’s ML pipeline stages (e.g., model.py, train.py, or preprocess.py) (amershi2019software).
@@
    Code check: If the previous steps did not yield valuable information, we conducted a finer-grained examination, which entailed 1) checking for the presence of ML library imports (e.g., tensorflow, sklearn, keras) and 2) inspecting for ML implementation specifics. The rationale for checking ML imports is based on the fact that ML-based systems utilize these libraries to build ML pipelines.
@@
4.2. Selection of Non-ML Projects
To answer the RQ1 and RQ4, we also need a set of non-ML projects as a baseline. To allow a fair comparison, we employ a consistent selection process for both ML and non-ML projects.
Specifically, we followed the sampling process outlined in Subsection 4.1 and collected 318 non-ML Python projects from GitHub (matching the count of ML projects). The C1, C2, C3, and C5 ML criteria are directly reused for the non-ML set, whereas, the C4 criterion is adapted so that, while the ML criterion specifies that repositories must implement ML, the non-ML criterion explicitly excludes ML implementations.
Using the same keyword-based selection mechanism of ML projects, we used the keywords “python projects”, “server” and “” (empty keywords) to obtain a list of non-ML python projects sorted by the “best match” criterion of GitHub Search. The rationale for using “server” is that this keyword often indicates backend implementations and is a non-specific term, i.e., the term “server” can have a diverse meaning in different software contexts, making it not specific to a technology, but a general term suitable for identifying non-ML repositories. Such projects often entail complex architectural decisions, configuration management challenges, and performance optimization issues, which are fertile grounds for the accumulation of technical debt.
@@ 4.4. SATD Detector Tool
As stated in our literature review Section 2.3, many SATD detector tools used to identify instances of SATD from code comments. For our study, we use the state of the art SATD detection tool [^7] Accessed on: 2024-03-22, https://github.com/Tbabm/SATDDetector-Core, proposed by Liu et al. (liu2018satd), which employs a NLP-based pre-trained text classification ML model to automatically detect the presence of SATD in a code comment. As compared to other contemporary SATD detection tools that rely on pattern matching, this tool demonstrates a higher recall (liu2018satd).
This tool’s use of an NLP-based model to determine whether a specific comment signifies SATD contrasts most of the other approaches, which merely scan for words like “todo” or “fixme” using string matching techniques. Therefore, the tool can successfully identify regular comments such as “300 iterations seems good enough but you can certainly train longer” or “naive approach for generating forward mapping this is naive and probably not robust” as SATD even when such strings lack an explicitly defined “TODO”. This SATD detector has been utilized in previous studies. For instance, Liu et al. (liu2021exploratory) used this tool for studying SATD in deep learning frameworks, while Zampetti et al. (Zampetti) used this tool for recommending when design technical debt should be self admitted.